Find the directional derivative of  at point  in the direction of vector. Where

 ,  and

F(x,y,z)=z-x/z+y, p (3,4,-5) and u= 2i+6j-k



1
Expert's answer
2021-02-28T17:08:15-0500

Directional derivative of function "F(x,y,z)=\dfrac{z-x}{z+y}" along direction of vector "\bold{u}=2\bold{i}+6\bold{j}-\bold{k}" at the point "P(3,4,-5)" is


"\dfrac{\partial F(P)}{\partial\bold{u}}=" "\dfrac{\partial F(P)}{\partial x}\cos\alpha+\dfrac{\partial F(P)}{\partial y}\cos\beta+\dfrac{\partial F(P)}{\partial z}\cos\gamma" ,


where

"\dfrac{\partial F}{\partial x}=" "\dfrac{\partial}{\partial x}\left(\dfrac{z-x}{z+y}\right)=" "\dfrac{1}{z+y}\cdot\dfrac{\partial}{\partial x}\left(z-x\right)=" "\dfrac{1}{z+y}\cdot\left(-1\right)=" "-\dfrac{1}{z+y}" ,


"\dfrac{\partial F(P)}{\partial x}=-\dfrac{1}{-5+4}=-\dfrac{1}{-1}=1" ,


"\dfrac{\partial F}{\partial y}=" "\dfrac{\partial}{\partial y}\left(\dfrac{z-x}{z+y}\right)=" "(z-x)\cdot\dfrac{\partial}{\partial y}\left(\dfrac{1}{z+y}\right)=" "(z-x)\cdot\left(-\dfrac{1}{(z+y)^2}\right)="


"-\dfrac{z-x}{(z+y)^2}" ,


"\dfrac{\partial F(P)}{\partial y}=-\dfrac{-5-3}{(-5+4)^2}=-\dfrac{-8}{(-1)^2}=\dfrac{8}{1}=8" ,


"\dfrac{\partial F}{\partial z}=" "\dfrac{\partial}{\partial z}\left(\dfrac{z-x}{z+y}\right)=" "\dfrac{\left(\dfrac{\partial}{\partial z}(z-x)\right)\cdot(z+y)-(z-x)\cdot\dfrac{\partial}{\partial z}(z+y)}{(z+y)^2}="


"\dfrac{1\cdot(z+y)-(z-x)\cdot1}{(z+y)^2}=" "\dfrac{z+y-z+x}{(z+y)^2}=\dfrac{y+x}{(z+y)^2}" ,


"\dfrac{\partial F(P)}{\partial z}=" "\dfrac{4+3}{(-5+4)^2}=\dfrac{7}{(-1)^2}=\dfrac{7}{1}=7" ,


"|\bold{u}|=\sqrt{2^2+6^2+(-1)^2}=\sqrt{4+36+1}=\sqrt{41}"


"\cos\alpha=\dfrac{2}{\sqrt{41}}" , "\cos\beta=\dfrac{6}{\sqrt{41}}" , "\cos\gamma=-\dfrac{1}{\sqrt{41}}" .


So,


"\dfrac{\partial F(P)}{\partial\bold{u}}=1\cdot\dfrac{2}{\sqrt{41}}+8\cdot\dfrac{6}{\sqrt{41}}-7\cdot\dfrac{1}{\sqrt{41}}=" "\dfrac{2+48-7}{\sqrt{41}}=\dfrac{43}{\sqrt{41}}" .